Hello I need help for this. Thank you.



1. A balloon has a volume of 4.0 L when at sea level (1.0 atm) at a room temperature of 28 °C . What will be its volume when inflated with the same amount of gas at an elevation where the atmospheric pressure is 700 mm Hg at 28 °C?

2. A gas sample occupies a volume of 12.0 L at 25 °C and 750 torr. At what temperature in °C will the gas occupy a volume of 15.0 L at 750 torr?

3. The volume of a gas sample at 0°C and 1.0 atm is 10.0 L. How many moles of gas are contained in the sample?

4. What will be the value of the universal gas constant if pressure is in kPa and volume is in liters?


1
Expert's answer
2021-01-19T04:37:27-0500

1. pV=nRT. n=pV/RT = 1 x 4 / 0.082 x 301 = 0,162061421.

V= nRT/p= 0,162061421 x 62400 x 301/700 = 4 348,43206 ml.

2. 750 torr= 100 koa.

pV=nRT. n= pV/RT = 100 x 12/8.314 x 298 = 0.4843 mol.

T=pV/nR = 100 x 15 / 0.4843 x 8.314 =

372,534733 K.

3. n=pV/RT = 1 x 10/ 0.082 x 273 = 0,446707766 moles.

4. 8.314 J/mol x K.